OpenStudy (anonymous):

find the derivative of (secx+cscx)/cscx Need step by step I get lost some where just don't know where

OpenStudy (debbieg):

I would suggest that you start by re-writing the expression, using trig identities: \[\frac{ \sec(x)+\csc(x) }{\csc(x) }=\frac{ \sec(x) }{\csc(x) }+\frac{ \csc(x) }{\csc(x) }=\tan(x)+1\] Do you see why this is so? Then you just need a couple of simple derivative rules. You probably have a list of them somewhere, but you can check here if not: http://www.mathsisfun.com/calculus/derivatives-rules.html

OpenStudy (anonymous):

Sorry but the correct answer is sec^2x

OpenStudy (anonymous):

nevermind I see what you are telling me

OpenStudy (debbieg):

Yes, that's what I got too. :)

OpenStudy (anonymous):

Thank You!!!! A lot easier to do it that way!!!! didn't see it!